# Dissipative Yao-Lee Spin-Orbital Model: Exact Solvability and $\mathcal{P}\mathcal{T}$ Symmetry Breaking

### v1

@@ Abstract
Exactly solvable dissipative models provide an analytical tool for studying the relaxation dynamics in open quantum systems. In this work, we study an exactly solvable model based on an anisotropic variant of the Yao-Lee spin-orbital model, with dissipation acting in the spin sector. We map Liouvillian dynamics to fermions hopping in a doubled Hilbert space under a non-Hermitian Hamiltonian, and demonstrate the model’s exact solvability. We analyze the strong and weak symmetries of the model, which protect an exponentially large manifold of non-equilibrium steady states, establishing the system as a physically feasible dissipative spin liquid. Furthermore, we analyze the transient dynamics in a translationally invariant sector and discover that the single-particle Liouvillian spectrum hosts an exceptional ring in momentum space. We map out a characteristic $\mathcal{P}\mathcal{T}$ symmetry breaking transition driven by the dissipation strength, which governs the crossover from oscillatory to decaying relaxation of physical observables. Our work provides a physically motivated, solvable setting for exploring the coexistence of dissipative spin liquid physics and non-Hermitian spectral topology.
@@ I Introduction
The non-equilibrium dynamics of quantum many-body systems coupled to an environment, through decoherence, dissipation, or measurements, has emerged as a rapidly growing frontier in condensed matter physics [open1, open2, open3, measurement3, weakmeasurement1, open4, open5, measurement2]. While the coupling to an external environment has traditionally been viewed as a source of detrimental noise, recent developments have established dissipation as a powerful resource, particularly with advances in experimental control across atomic, optical, solid-state, and superconducting platforms [np_Diehl2008, np_DallaTorre2010, np_Barreiro2011, np_Eisert2015, np_Skou2021, Özdemir2019, IBMQ]. These developments have stimulated substantial interest in understanding how system-environment couplings can be engineered to shape relaxation dynamics, steady-state structure, and realize topological phenomena in driven-dissipative settings [cenke_spinliquid, decoherence_lee, driven1, driven2, driven3, driven4, dissipativeengineering5, dissipativeengineering6, dissipativeengineer2, dissipativeengineer3, dissipativeengineer4].
@@
Despite recent progress, the theoretical study of open many-body quantum systems remains challenging due the exponential explosion of the operator space. Compared to closed quantum systems, where the Hilbert space dimension scales as $2^{L}$ for a system with $L$ qubits, open quantum systems require treating density matrices whose dimension scales as $4^{L}$. This scaling quickly renders exact numerical simulations unfeasible beyond small system sizes. Tensor-network-based approaches [TN_PhysRevLett.93.207204, TN_PhysRevLett.93.207205, TN_PhysRevLett.114.220601, TN_PhysRevLett.116.237201, TN_PhysRevLett.132.200403, TN_Jaschke_2019, TN_hmbj-lg7p] and quantum-trajectory methods [traj_Daley04032014, traj_PhysRevA.45.4879, traj_PhysRevLett.68.580, traj_RevModPhys.70.101, traj_xue2025simulationbilayerhamiltoniansbased] have extended our reach, but they remain numerically costly and may obscure the underlying physics. In contrast, exactly solvable Liouvillians [BCS_solvable, Shackleton_dissipative, arovas_dissipative, lucas_sa, exact1, exact2, exact3, exact4, exact5, exact6, exact7] provide non-perturbative insight into the complex behavior of relaxation and steady states, and constitute valuable benchmarks for new numerical techniques.
@@
Recently, a notable development along this line of inquiry is the dissipative spin liquids (DSL): exactly solvable Liouvillians with an extensive number of symmetries, which protect an exponentially large manifold of non-equilibrium steady states [Shackleton_dissipative, arovas_dissipative, lucas_sa]. However, existing DSL constructions are formulated in terms of Gamma matrices acting on enlarged on-site Hilbert spaces and lack a direct mapping to local spin degrees of freedom. This makes it difficult to assess how such models might be connected to realistic materials or solid-state platforms.
@@
In this work, we introduce an exactly solvable dissipative model that is both physically motivated and dynamically rich. We consider an anisotropic variant of the Yao-Lee spin-orbital model [yaolee_original], a well-studied generalization of Kitaev’s honeycomb model [Kitaev_honeycomb] that incorporates orbital (pseudospin) degrees of freedom while retaining its exact solvability. The Yao-Lee model hosts a variety of interesting phenomena, such as spin-1 fermionic excitations, vison crystallization, and topological band transitions in the $\sigma$ sector [yaolee_original, yaolee1, yaolee2, yaolee3]. More recently, Ref. [yaolee_experiment] presents a road-map to realizing the Yao-Lee model in experiments. While non-Hermitian extension of the model has been studied [nonhermitian_yaolee], how Lindbladian dissipation affects its spin-orbital dynamics and steady state structure has not been explored.
@@
We solve the dissipative model through third quantization [third_quantization], mapping the Liouvillian super-operator to a non-Hermitian Hamiltonian that is quadratic in fermions hopping on a bilayer honeycomb lattice. This mapping allows for an analysis of the system’s strong and weak symmetries, from which we obtain a systematic criterion for determining flux sectors that contain non-equilibrium steady states (NESS). We identify an exponentially large manifold of NESS, which confirms that the model realizes a DSL [Shackleton_dissipative, arovas_dissipative, lucas_sa], but in a setting where the local degrees of freedom are more closely connected to experiments.
@@
While the NESS manifold characterizes the long-time limit, the relaxation to equilibrium contains rich transient dynamics. To explore this, we focus on a translationally invariant flux sector and compute the single-particle Liouvillian spectrum analytically. We find that the single-particle spectrum hosts an exceptional ring in momentum space: a continuous contour of exceptional points (EPs), where eigenvalues and eigenvectors coalesce [Huber_PT, prosen_PT, Nakanishi_PT, dissipative_stark, quantum_quench_drivendissipative, localizer].
@@ II Mapping between Lindbladians and bilayer Hamiltonians
We first briefly review Lindbladian dynamics and the mapping from Lindbladians to non-Hermitian Hamiltonians on a doubled Hilbert space. Given a Hamiltonian $H$ and a set of Markovian jump operators $\{L_{i}\}$, the density matrix $\rho$ of an open quantum system evolves according to the Lindblad master equation [lindblad, Gorini]:
@@
The Liouvillian $\mathcal{L}$ is a completely positive, trace-preserving (CPTP) super-operator, mapping between bounded operators on the original Hilbert space i.e. $\mathcal{L}$ is a map from $B(\mathcal{H})\rightarrow B(\mathcal{H})$. If we treat the space of operators $B(\mathcal{H})$ as a Hilbert space by itself, equipped with the inner product $\langle\rho,\sigma\rangle=\operatorname{Tr}(\rho^{\dagger}\sigma)$, then $\mathcal{L}$ can be viewed as an operator acting on this space.
In this representation, the density matrix $\rho$ is vectorized by the Choi–Jamiołkowski isomorphism [CHOI1975285, JAMIOLKOWSKI1972275]:
@@
Beyond the correspondence between spectra and steady states, the mapping is especially useful for organizing the symmetry structure of Liouvillians. The doubled-Hilbert-space representation provides a natural framework to classify and analyze symmetries of the Liouvillian [Shinsei_symmetry, L_symmetry, L_symmetry2, L_symmetry3, prosen_PT]. For example, breaking of parity–time ($\mathcal{P}\mathcal{T}$) symmetry in the Liouvillian [prosen_PT, Huber_2020, PT_open] appears as the emergence of exceptional points (EPs) in the non-Hermitian Hamiltonian, where eigenvalues and eigenvectors coalesce [El-Ganainy2018, Özdemir2019, RevModPhys.93.015005, PhysRevResearch.4.L042025, Ding_2022]. In Sec. IV, we detail how symmetries of our model are interpreted in the doubled-Hilbert-space formalism.
Figure 1: (a) Illustration of the mapping between an open system described by the Liouvillian $\mathcal{L}$ (left) and the non-Hermitian bilayer Hamiltonian $\mathcal{H}$ (right). In the doubled Hilbert space, the jump operators couple the bra and ket spaces. The two sublattices $A$ and $B$ are labeled by blue and red dots, respectively. (b) A hexagonal plaquette in the honeycomb lattice. There are three types of bonds: $x$, $y$, and $z$. (c) The Brillouin zone of the honeycomb lattice.
@@ III Exactly solvable dissipative spin orbital model
Our model is a variant of the Yao-Lee (YL) model [yaolee_original], which generalizes Kitaev’s honeycomb model [Kitaev_honeycomb] to a spin-orbital setting. The YL Hamiltonian reads:
@@
where $N$ is the number of sites in one layer. Throughout this paper, the overhead tilde will denote operators that act on the ket copy of the Hilbert space, $\tilde{A}:=\mathbb{I}\otimes A^{T}$ [CHOI1975285, JAMIOLKOWSKI1972275].
To solve this model $\mathcal{H}$, we introduce two types of Majorana fermions $c$ and $d$ via the transformation that has been used in the solution of the Yao-Lee model [yaolee_original]:
@@
The Majorana fermions satisfy mutual anticommutation relations, $\{c_{i}^{\alpha},c_{j}^{\beta}\}=\{d_{i}^{\alpha},d_{j}^{\beta}\}=\delta_{ij}^{\alpha\beta}$, $\{c_{i}^{\alpha},d_{j}^{\beta}\}=0$. Similar decompositions are performed on $\tilde{\sigma}$ and $\tilde{\tau}$. Compared to other exactly solvable dissipative models involving Gamma matrices [arovas_dissipative, Shackleton_dissipative, lucas_sa], the Majoranas in this model has a more direct connection to the physics variables: namely, the $c$ and $d$ Majoranas are associated with the spin ($\sigma$) and orbital ($\tau$) degrees of freedom, respectively.
Since above decomposition into Majoranas (Eq. (9)) is over-complete [Kitaev_honeycomb], we need to project back to the physical space by imposing the local constraint:
@@
where $\Lambda_{i}$ and $\tilde{\Lambda}_{i}$ are $\pm 1$. The gauge-invariant quantities are therefore plaquette fluxes i.e. products of bond operator around plaquettes [Kitaev_honeycomb, yaolee_original]; see Sec. IV for a more detailed discussion.
@@ IV.1 Strong and Weak Symmetry
In open quantum systems, symmetries are defined with respect to the Liouvillian $\mathcal{L}$, and the above notion of symmetry can be generalized. There are two types of symmetries: strong and weak symmetry [prosen_note, PhysRevA.89.022118]. The interplay between these two types of symmetries, such as spontaneous strong-to-weak symmetry breaking, has drawn lots of recent interest [SSWSB_wangzhong, SSWSB2, SSWSB3, SWSSB_Lee2025symmetryprotected, SWSSB_PRXQuantum.4.030317, SWSSB_1xzd-g9s5].
@@
for some operator $S$. Note that $S,L_{i}]=[S,H]=0$ implies $\mathcal{S},\mathcal{L}]=0$, but not vice versa; a strong symmetry therefore necessarily implies a weak symmetry, but the converse is not true. While both types of symmetries allow for block-diagonalizing the Liouvillian, a strong symmetry has an associated conserved charge, while weak symmetries do not [prosen_note].
One can also understand the difference between strong and weak symmetries in the bilayer picture, in which $\mathcal{L}$ is mapped to a non-Hermitian Hamiltonian $\mathcal{H}$ in the doubled space. Both types of symmetries leave $\mathcal{H}$ invariant. However, weak symmetries must act jointly on both copies of the doubled Hilbert space to preserve $\mathcal{H}$, whereas strong symmetries only act on either copy independently [Ryu_classification]. Consequently, charges associated with strong symmetries are conserved on each layer independently. In contrast, weak symmetry only implies conservation of the difference between charges on the two layers, which do not translate to conserved observables of the original single-layer system [SSWSB_wangzhong].
The dissipative spin-orbital model we study has an extensive number of strong and weak symmetries. Indeed, similar to Kitaev’s honeycomb model [Kitaev_honeycomb], it is precisely the model’s extensive symmetries that allow for its exact solvability. To understand the strong symmetries, we first consider the dissipation-less limit, $\gamma=0$. The model reduces to a variant of the Yao-Lee model [yaolee_original], whose $\text{SU}(2)$ symmetry is explicitly broken. In this (closed) system, the conserved fluxes are products of bond operators around each hexagon [yaolee_original]:
@@
Next, we turn to the model’s weak symmetries. Following Refs. [Shackleton_dissipative, arovas_dissipative, lucas_sa], we consider the vertical plaquettes $V_{p}$ that link the two layers, as illustrated in Fig. 2. Let $i$ and $j$ denote two nearest neighbors on the original layer, connected by a bond of type $\lambda\in\{x,y,z\}$. We define:
@@
Compared to previously studied exactly solvable DSL models [arovas_dissipative, Shackleton_dissipative, lucas_sa], our system hosts an additional, weak $\text{U}(1)$ symmetry under the following super-operator:
@@ IV.2 Steady States
Unlike closed (Hermitian) systems, where Lieb’s Theorem [lieb_theorem] dictates that the ground state lies in the sector with zero flux, there are no generic principles for choosing flux configurations in open quantum systems. Nevertheless, strong and weak symmetries still provide useful guidelines for counting non-equilibrium steady states (NESS) of the system.
@@
Beyond the two limits, we need to analyze using symmetries. As shown in Ref. [prosen_note], the number of NESS is at least the number of distinct eigenvalues of strong symmetries in the system. Indeed, the Liouvillian is block-diagonalized under left and right actions by strong symmetries, and steady states live in the “diagonal” blocks where the left and right labels coincide. In the bilayer representation, this enforces matched hexagonal fluxes, $W_{p}=\widetilde{W}_{p}$, for every plaquette $p$, so each steady sector is labeled by a common flux pattern across both layers. Note that each of the hexagonal plaquettes $W_{p}$ can be +1+1 or 1-1; the only requirement is a matched pattern on the two layers.
Next we analyze the weak symmetries. Following Ref. [prosen_note], we will assume that each sector labeled by eigenvalues of the weak symmetries contains at most one NESS. This behavior is typical in local dissipative models [prosen_note]. Under this assumption, a sector can host an NESS only if it transforms trivially under all weak symmetries (i.e. all weak symmetries have eigenvalue 1). To see this, let $\mathcal{U}$ be a weak symmetry acting as $\mathcal{U}\rho=V\rho V^{\dagger}$, where $V$ is a unitary operator. If $\rho_{\text{ss}}$ is the unique steady state that satisfies $\mathcal{L}\rho_{\text{ss}}=0$, then:
@@
By uniqueness, $\mathcal{U}\rho_{ss}=u\rho_{ss}$ for some eigenvalue $u$. Taking the trace on both sides yields $\operatorname{Tr}(V\rho_{ss}V^{\dagger})=\operatorname{Tr}(\rho_{ss})=u\operatorname{Tr}(\rho_{ss})\rightarrow u=1$. For our model, these constraints select weak-symmetry sectors with a uniform interlayer flux parity $V_{p}=-1$ on every plaquette. Note that an eigenvalue of $1$ under the weak symmetry in Eq. (26) implies $V_{p}=-1$ due to the minus sign in the definition of $V_{p}$ in Eq. (25). It also constraints steady states to sectors with zero relative U(1)(1) charge, $\sum_{i}\sigma_{i}^{y}-\tilde{\sigma}_{i}^{y}=0$. In terms of the fermionic variables, the requirement for zero relative U(1)(1) translates to:
@@
We can count the number of NESS following the above analysis. On a honeycomb lattice with $N$ sites and periodic boundary conditions, there are $E=\frac{3}{2}N$ edges. Using Euler’s formula for planar graph V+EF=2V+E-F=2, where $V$, $E$, $F$ denote the number of vertices, edges, and faces respectively, we see that there are $N/2$ hexagonal plaquettes. The plaquette fluxes satisfy the global constraint $\prod_{p}W_{p}=1$, leaving $\frac{N}{2}-1$ independent fluxes. In addition, there are two independent, non-contractible $\mathbb{Z}_{2}$ Wilson loops. Hence the number of diagonal sectors under decomposition under the strong symmetries is $2^{N/2-1}\times 2^{2}=2^{N/2+1}$. Imposing the weak-symmetry constraints yields exactly one NESS per (weak symmetry) sector. Therefore, the number of NESS is $2^{N/2+1}$; this may reduce to $2^{N/2-1}$ if global Wilson loops are pinned by boundary conditions. The resulting exponentially large number of NESS demonstrates that our model realizes a type of “dissipative spin liquid”, as discussed in Refs. [arovas_dissipative, lucas_sa, Shackleton_dissipative].

### v2

@@ Abstract
Exactly solvable dissipative models provide an analytical tool for studying the relaxation dynamics in open quantum systems. In this work, we study an exactly solvable model based on an anisotropic variant of the Yao-Lee spin-orbital model, with dissipation acting in the spin sector. We map Liouvillian dynamics to fermions hopping in a doubled Hilbert space under a non-Hermitian Hamiltonian and demonstrate the model’s exact solvability. We analyze the model’s strong and weak symmetries, which protect an exponentially large manifold of non-equilibrium steady states, establishing the system as a physically feasible dissipative spin liquid. Furthermore, we analyze the transient dynamics in a translationally invariant sector and discover that the single-particle Liouvillian spectrum hosts an exceptional ring in momentum space. We map out a characteristic $\mathcal{P}\mathcal{T}$ symmetry breaking transition driven by the dissipation strength, which governs the crossover from oscillatory to decaying relaxation of physical observables. Our work provides a physically motivated, solvable setting for exploring the coexistence of dissipative spin liquid physics and Liouvillian spectral singularities.
@@ I Introduction
The non-equilibrium dynamics of quantum many-body systems coupled to an environment, through decoherence, dissipation, or measurements, has emerged as a rapidly growing frontier in condensed matter physics [67, 68, 9, 39, 79, 37, 26, 87]. While the coupling to an external environment has traditionally been viewed as a source of detrimental noise, recent developments have established dissipation as a powerful resource, particularly with advances in experimental control across atomic, optical, solid-state, and superconducting platforms [21, 20, 5, 24, 77, Özdemir2019, 28]. These developments have stimulated substantial interest in understanding how system-environment couplings can be engineered to shape relaxation dynamics, steady-state structure, and realize topological phenomena in driven-dissipative settings [80, 40, 65, 49, 58, 15, 43, 84, 42, 72, 11].
@@
Despite recent progress, the theoretical study of open many-body quantum systems remains challenging due the exponential explosion of the operator space. Compared to closed quantum systems, where the Hilbert space dimension scales as $2^{L}$ for a system with $L$ qubits, open quantum systems require treating density matrices whose dimension scales as $4^{L}$. This scaling quickly renders exact numerical simulations unfeasible beyond small system sizes. Tensor-network-based approaches [81, 90, 16, 82, 47, 34, 57] and quantum-trajectory methods [18, 23, 19, 59, 83] have extended our reach, but they remain numerically costly and may obscure the underlying physics. In contrast, exactly solvable Liouvillians [17, 73, 29, 69, 54, 50, 14, 76, 75, 51, 66] provide non-perturbative insight into the complex behavior of relaxation and steady states, and constitute valuable benchmarks for new numerical techniques.
@@
Recently, a notable development along this line of inquiry is the dissipative spin liquids (DSL): exactly solvable Liouvillians with an extensive number of symmetries, which protect an exponentially large manifold of non-equilibrium steady states [73, 29, 69]. However, existing DSL constructions are formulated in terms of Gamma matrices acting on enlarged on-site Hilbert spaces and lack a direct mapping to local spin degrees of freedom. This makes it difficult to assess how such models might be connected to realistic materials or solid-state platforms.
@@
In this work, we introduce an exactly solvable dissipative model that is both physically motivated and dynamically rich. We consider an anisotropic variant of the Yao-Lee spin-orbital model [88], a well-studied generalization of Kitaev’s honeycomb model [36] that incorporates orbital (pseudospin) degrees of freedom while retaining its exact solvability. The Yao-Lee model hosts a variety of interesting phenomena, such as spin-1 fermionic excitations, vison crystallization, and topological band transitions in the $\sigma$ sector [88, 2, 60, 1]. More recently, Ref. [13] presents a road-map to realizing the Yao-Lee model in experiments. While non-Hermitian extension of the model has been studied [48], how Lindbladian dissipation affects its spin-orbital dynamics and steady state structure has not been explored.
@@
We solve the dissipative model through third quantization [61], mapping the Liouvillian super-operator to a non-Hermitian Hamiltonian that is quadratic in fermions hopping on a bilayer honeycomb lattice. This mapping allows for an analysis of the system’s strong and weak symmetries, from which we obtain a systematic criterion for determining flux sectors that contain non-equilibrium steady states (NESS). We identify an exponentially large manifold of NESS, which confirms that the model realizes a DSL [73, 29, 69], but in a setting where the local degrees of freedom are more closely connected to experiments.
@@
While the NESS manifold characterizes the long-time limit, the relaxation to equilibrium contains rich transient dynamics. To explore this, we focus on a translationally invariant flux sector and compute the single-particle Liouvillian spectrum analytically. We find that the single-particle spectrum hosts an exceptional ring in momentum space: a continuous contour of exceptional points (EPs), where eigenvalues and eigenvectors coalesce [32, 63, 55, 89, 78, 64].
@@ II Mapping between Lindbladians and bilayer Hamiltonians
We first briefly review Lindbladian dynamics and the mapping from Lindbladians to non-Hermitian Hamiltonians on a doubled Hilbert space. Given a Hamiltonian $H$ and a set of Markovian jump operators $\{L_{i}\}$, the density matrix $\rho$ of an open quantum system evolves according to the Lindblad master equation [46, 30]:
@@
The Liouvillian $\mathcal{L}$ is the generator of a quantum dynamical semigroup, mapping between bounded operators on the original Hilbert space, $\mathcal{L}:B(\mathcal{H})\to B(\mathcal{H})$. If we treat the space of operators $B(\mathcal{H})$ as a Hilbert space by itself, equipped with the inner product $\langle\rho,\sigma\rangle=\operatorname{Tr}(\rho^{\dagger}\sigma)$, then $\mathcal{L}$ can be viewed as an operator acting on this space.
In this representation, the density matrix $\rho$ is vectorized by the Choi–Jamiołkowski isomorphism [12, 33]:
@@
Beyond the correspondence between spectra and steady states, the mapping is especially useful for organizing the symmetry structure of Liouvillians. The doubled-Hilbert-space representation provides a natural framework to classify and analyze symmetries of the Liouvillian [35, 70, 45, 4, 63]. For example, breaking of parity–time ($\mathcal{P}\mathcal{T}$) symmetry in the Liouvillian [63, 32, 56] appears as the emergence of exceptional points (EPs) in the non-Hermitian Hamiltonian, where eigenvalues and eigenvectors coalesce [25, Özdemir2019, 8, 86, 22]. In Sec. IV, we detail how symmetries of our model are interpreted in the doubled-Hilbert-space formalism.
Figure 1: (a) Illustration of the mapping between an open system described by the Liouvillian $\mathcal{L}$ (left) and the non-Hermitian bilayer Hamiltonian $\mathcal{H}$ (right). In the doubled Hilbert space, the jump operators couple the bra and ket spaces. The two sublattices AA and $B$ are labeled by blue and red dots, respectively. (b) A hexagonal plaquette in the honeycomb lattice. There are three types of bonds: $x$, $y$, and $z$. (c) The Brillouin zone of the honeycomb lattice.
@@ III Exactly solvable dissipative spin orbital model
Our model is a variant of the Yao-Lee (YL) model [88], which generalizes Kitaev’s honeycomb model [36] to a spin-orbital setting. The YL Hamiltonian reads:
@@
where NN is the number of sites in one layer. Throughout this paper, the overhead tilde will denote operators that act on the ket copy of the Hilbert space, $\tilde{A}:=\mathbb{I}\otimes A^{T}$ [12, 33].
To solve this model $\mathcal{H}$, we introduce two types of Majorana fermions $c$ and $d$ via the transformation that has been used in the solution of the Yao-Lee model [88]:
@@
The Majorana fermions satisfy mutual anticommutation relations, $\{c_{i}^{\alpha},c_{j}^{\beta}\}=\{d_{i}^{\alpha},d_{j}^{\beta}\}=\delta_{ij}^{\alpha\beta}$, $\{c_{i}^{\alpha},d_{j}^{\beta}\}=0$. Similar decompositions are performed on $\tilde{\sigma}$ and $\tilde{\tau}$. Compared to other exactly solvable dissipative models involving Gamma matrices [29, 73, 69], the Majoranas in this model has a more direct connection to the physics variables: namely, the $c$ and $d$ Majoranas are associated with the spin ($\sigma$) and orbital ($\tau$) degrees of freedom, respectively.
Since above decomposition into Majoranas (Eq. (9)) is over-complete [36], we need to project back to the physical space by imposing the local constraint:
@@
where $\Lambda_{i}$ and $\tilde{\Lambda}_{i}$ are $\pm 1$. The gauge-invariant quantities are therefore plaquette fluxes i.e. products of bond operator around plaquettes [36, 88]; see Sec. IV for a more detailed discussion.
@@ IV.1 Strong and Weak Symmetry
In open quantum systems, symmetries are defined with respect to the Liouvillian $\mathcal{L}$, and the above notion of symmetry can be generalized. There are two types of symmetries: strong and weak symmetry [10, 3]. The interplay between these two types of symmetries, such as spontaneous strong-to-weak symmetry breaking, has drawn lots of recent interest [31, 41, 71, 40, 39, 27].
@@
for some operator $S$. Note that $S,L_{i}]=[S,H]=0$ implies $\mathcal{S},\mathcal{L}]=0$, but not vice versa; a strong symmetry therefore necessarily implies a weak symmetry, but the converse is not true. While both types of symmetries allow for block-diagonalizing the Liouvillian, a strong symmetry has an associated conserved charge, while weak symmetries do not [10].
One can also understand the difference between strong and weak symmetries in the bilayer picture, in which $\mathcal{L}$ is mapped to a non-Hermitian Hamiltonian $\mathcal{H}$ in the doubled space. Both types of symmetries leave $\mathcal{H}$ invariant. However, weak symmetries must act jointly on both copies of the doubled Hilbert space to preserve $\mathcal{H}$, whereas strong symmetries only act on either copy independently [35]. Consequently, charges associated with strong symmetries are conserved on each layer independently. In contrast, weak symmetry only implies conservation of the difference between charges on the two layers, which do not translate to conserved observables of the original single-layer system [31].
The dissipative spin-orbital model we study has an extensive number of strong and weak symmetries. Indeed, similar to Kitaev’s honeycomb model [36], it is precisely the model’s extensive symmetries that allow for its exact solvability. To understand the strong symmetries, we first consider the dissipation-less limit, $\gamma=0$. The model reduces to a variant of the Yao-Lee model [88], whose $\text{SU}(2)$ symmetry is explicitly broken. In this (closed) system, the conserved fluxes are products of bond operators around each hexagon [88]:
@@
Next, we turn to the model’s weak symmetries. Following Refs. [73, 29, 69], we consider the vertical plaquettes $V_{p}$ that link the two layers, as illustrated in Fig. 2. Let $i$ and $j$ denote two nearest neighbors on the original layer, connected by a bond of type $\lambda\in\{x,y,z\}$. We define:
@@
Compared to previously studied exactly solvable DSL models [29, 73, 69], our system hosts an additional, weak $\text{U}(1)$ symmetry under the following super-operator:
@@ IV.2 Steady States
Unlike closed (Hermitian) systems, where Lieb’s Theorem [44] dictates that the ground state lies in the sector with zero flux, there are no generic principles for choosing flux configurations in open quantum systems. Nevertheless, strong and weak symmetries still provide useful guidelines for counting non-equilibrium steady states (NESS) of the system.
@@
Beyond the two limits, we need to analyze using symmetries. As shown in Ref. [10], the number of NESS is at least the number of distinct eigenvalues of strong symmetries in the system. Indeed, the Liouvillian is block-diagonalized under left and right actions by strong symmetries, and steady states live in the “diagonal” blocks where the left and right labels coincide. In the bilayer representation, this enforces matched hexagonal fluxes, $W_{p}=\widetilde{W}_{p}$, for every plaquette $p$, so each steady sector is labeled by a common flux pattern across both layers. Note that each of the hexagonal plaquettes $W_{p}$ can be +1+1 or 1-1; the only requirement is a matched pattern on the two layers.
Next we analyze the weak symmetries. Following Ref. [10], we will assume that each sector labeled by eigenvalues of the weak symmetries contains at most one NESS. This behavior is typical in local dissipative models [10]. Under this assumption, a sector can host an NESS only if it transforms trivially under all weak symmetries (i.e. all weak symmetries have eigenvalue 1). To see this, let $\mathcal{U}$ be a weak symmetry acting as $\mathcal{U}\rho=V\rho V^{\dagger}$, where $V$ is a unitary operator. If $\rho_{\text{ss}}$ is the unique steady state that satisfies $\mathcal{L}\rho_{\text{ss}}=0$, then:
@@
By uniqueness, $\mathcal{U}\rho_{ss}=u\rho_{ss}$ for some eigenvalue $u$. Taking the trace on both sides yields $\operatorname{Tr}(V\rho_{ss}V^{\dagger})=\operatorname{Tr}(\rho_{ss})=u\operatorname{Tr}(\rho_{ss})\rightarrow u=1$. For our model, these constraints select weak-symmetry sectors with a uniform interlayer flux parity $V_{p}=-1$ on every plaquette. Note that an eigenvalue of 11 under the weak symmetry in Eq. (26) implies $V_{p}=-1$ due to the minus sign in the definition of $V_{p}$ in Eq. (25). It also constraints steady states to sectors with zero relative U(1)(1) charge, $\sum_{i}\sigma_{i}^{y}-\tilde{\sigma}_{i}^{y}=0$. In terms of the fermionic variables, the requirement for zero relative U(1)(1) translates to:
@@
We can count the number of NESS following the above analysis. On a honeycomb lattice with NN sites and periodic boundary conditions, there are $E=\frac{3}{2}N$ edges. Using Euler’s formula for planar graph V+EF=2V+E-F=2, where $V$, $E$, $F$ denote the number of vertices, edges, and faces respectively, we see that there are $N/2$ hexagonal plaquettes. The plaquette fluxes satisfy the global constraint $\prod_{p}W_{p}=1$, leaving $\frac{N}{2}-1$ independent fluxes. In addition, there are two independent, non-contractible $\mathbb{Z}_{2}$ Wilson loops. Hence the number of diagonal sectors under decomposition under the strong symmetries is $2^{N/2-1}\times 2^{2}=2^{N/2+1}$. Imposing the weak-symmetry constraints yields exactly one NESS per (weak symmetry) sector. Therefore, the number of NESS is $2^{N/2+1}$; this may reduce to $2^{N/2-1}$ if global Wilson loops are pinned by boundary conditions.
The exponentially large number of NESS establishes our model as an example of “dissipative spin liquid” [29, 69, 73]. Dissipative spin liquids are generalizations of quantum spin liquids to the dissipative setting, in which Lindbladian dynamics preserves an emergent gauge structure, enforced by an extensive set of strong symmetries. The macroscopically degenerate manifold of NESS supports nonlocal correlations in the steady states, providing an exactly solvable setting for studying mixed state order [69]. Furthermore, exactly solvable dissipative models admit fractionalized descriptions, where decay modes fall into distinct sectors with parametrically separated relaxation timescales [73]. Our dissipative Yao-Lee model therefore serves as a solvable, experimentally relevant playground for investigating these exotic non-equilibrium phenomena.